### v1

@@ Abstract
We introduce several novel probabilistic quantum algorithms that overcome the normal unitary restrictions in quantum machine learning by leveraging the Linear Combination of Unitaries (LCU) method. Among our contributions are quantum native implementations of Residual Networks (ResNet); demonstrating a path to avoiding barren plateaus while maintaining the complexity of models that are hard to simulate classically. Furthermore, by generalising to allow control of the strength of residual connections, we show that the lower bound of the LCU success probability can be set to any arbitrary desired value. We also implement a quantum analogue of average pooling layers from convolutional networks. Our empirical analysis demonstrates that the LCU success probability remains stable for the MNIST database, unlocking a potential quadratic advantage in terms of image size compared to classical techniques. Finally, we propose a general framework for irreducible subspace projections for quantum encoded data. Using this, we demonstrate a novel rotationally invariant encoding for point cloud data via Schur-Weyl duality. We also show how this framework can be used to parameterise and control the amount of symmetry in an encoding; demonstrating improved classification performance for partially permutation invariant encoded point cloud data when compared to non-invariant or fully permutation invariant encodings. These new general algorithmic frameworks are all constructed under the same LCU method, suggesting that even more novel algorithms could be achieved by utilising the LCU technique.
keywords:
Quantum Machine Learning, Residual Networks, Quantum ResNet, Encoding Methods, 3D Computer Vision, Quantum Convolutional Neural Networks, Geometric Quantum Machine Learning, Point Cloud Data, Non-unitary Quantum Machine Learning
1 Introduction
@@
Quantum computing is an emerging technology with the potential to solve certain problems far more efficiently than classical techniques Shor_1997 ; Cao_2019 ; herman2022survey . The search for useful applications of quantum computers within the domain of machine learning is of particular importance given the significant impact classical machine learning has had on many fields and industries in recent years touvron2023llama ; chang2023muse ; frieder2023mathematical . Various quantum machine learning algorithms have been proposed, with some of the most common examples attempting to recreate a quantum analogue of the classical neural network Abbas_2021 ; Beer_2020 . However, the inherently unitary operations of quantum algorithms impose significant limitations, as they restrict the implementation of non-unitary operations that are essential in many classical machine learning models. To address this issue, we explore the Linear Combination of Unitaries (LCU) method Childs2012LCU ; Berry_2015 , a technique that employs ancilla qubits to probabilistically implement a linear combination of unitary operations, which therefore permits the implementation of non-unitary operations in quantum circuits. This unlocks the possibility of implementing non-unitary operations within quantum machine learning (QML) applications. We utilise this to present several novel applications of the LCU method with various advantages in the field of quantum machine learning.
@@
In Section 2 we adapt the classical residual learning framework to quantum variational circuits, facilitating a quantum native implementation of Residual Networks (ResNet) he2015deep ; marion2022scaling . We show that this quantum ResNet may provide a method of avoiding barren plateaus in Variational Quantum Circuit (VQC) models while also retaining complexity in the models such that they are overall difficult to simulate classically. Furthermore, we show that by including terms that parameterise the strength of the residual connections, it is possible to increase the lower bound of the probability of success of the LCU procedure to any desired value. This helps to overcome one of the main challenges of LCU methods, the fact that they have a fixed probability of success each time they are implemented.
@@
In Section 3 we implement a quantum analogue of average pooling operations from convolutional neural networks (CNN) RawatCNN ; SHARMA2018377 ; SaffarCNN by using the LCU method, unlocking a new technique for future use in quantum convolutional neural networks. We find that the LCU probability of success will equal $1$ in the case where all pixels are of the same colour and will decrease when pixels in the local pooling window of dimension $D\times D$ become more diverse. For real-world images, we provide the intuition that the probability remains relatively stable since most pixels are similar locally, except at the edges of subjects in the image. This intuition is supported by empirical evidence on $N\times N$ pixel images in the MNIST deng2012mnist database, which shows that probability decreases but levels off to a finite value as $D$ increases, and shows no discernible trend when increasing image size $N$. This potentially offers a quadratic advantage in terms of $N$ compared to classical pooling layers, as quantum parallelism allows the averaging of all $N^{2}$ pixels simultaneously.
@@
In Section 4 we present a method for projecting quantum encoded data to any combination of irreducible representation subspaces of a finite group, presenting a general framework for implementing full or partial symmetries in the encoding step of quantum machine learning models. These techniques are intended to reduce the effective dimension of quantum encoded data in an effort to improve generalisation performance, which has been reported to decline as the number of qubits and hence the dimension of the encoded quantum states increases huang_power_2021 . We show that this technique can recreate previous work on permutation invariant encodings for point cloud data heredge2023permutation as a special case. Furthermore, we implement a novel rotationally invariant encoding for point cloud data using the new technique by leveraging Schur-Weyl duality. This results in an encoded quantum state for point cloud data that is invariant if classical input point cloud data is rotated in 3-dimensional space, hence strongly enforcing rotation invariance on any model that subsequently uses this rotationally invariant input state. We further show that any combination of projections can be implemented at once, allowing certain symmetric subspaces of the data to be amplified or contracted to give increased flexibility over the amount of symmetry in the encoding. We demonstrate that intermediate levels of permutation symmetry for point cloud encoded data leads to an improved classification performance when compared to a non-symmetric or a fully permutation symmetric encoding.
@@
Table 1: Summary of the algorithms introduced in this work using the LCU framework. $G$ is a finite compact group. $\lvert\psi_{r}\rangle$ is the portion of a state $\lvert\psi\rangle$ that occupies the subspace of the irreducible representation $r$, with $\lvert\psi_{\text{rot}}\rangle$ meaning the portion of $\lvert\psi\rangle$ that is in the rotationally invariant subspace. Note that $\langle\psi_{r}\rvert\psi_{r}\rangle$ is not necessarily equal to 1, as the condition is $\langle\psi\rvert\psi\rangle=\sum_{r}\langle\psi_{r}\rvert\psi_{r}\rangle$ = 1. The parameters $a_{r}$ can be chosen by the user subject to normalisation constraints. In average pooling we use a pooling window of dimension $D\times D$ pixels. $L$ is the number of layers in the quantum ResNet. $\beta_{l}\in[0,1],\beta_{l}\in\mathbb{R}$ parameterise the strength of the skipped connections in the quantum ResNet and can be freely adjusted. Throughout this work $\log\equiv\log_{2}$ is the logarithmic function in base 2.
Algorithm Motivation Ancilla Qubits Success Probability Section
@@
Quantum ResNet Residual Layers Potentially avoids barren plateaus in non-simulatable models $\mathcal{O}(L)$ $[\prod_{l}\big{(}1-4\beta_{l}(1-\beta_{l})\big{)},1]$ Sec 2
Average Pooling Potential quadratic speedup in $N$ over averaging classically $\mathcal{O}(\log(D))$ Shown for MNIST in Fig 7 and Fig 8. Sec 3
@@
Irreducible Subspace Projections Can freely enforce and parameterise symmetry of a quantum encoded state $\mathcal{O}(\log(\rvert G\rvert))$ $\sum_{r}\rvert a_{r}\rvert^{2}\langle\psi_{r}\rvert\psi_{r}\rangle$ Sec 4
@@
Rotationally Invariant Encoding Encoding respects rotational invariance of point cloud data $\mathcal{O}(\log(\rvert G\rvert))$ $\langle\psi_{\text{rot}}\rvert\psi_{\text{rot}}\rangle$ Sec 4.3.2
1.1 Quantum Native ResNet
@@
A common issue in the training of variational quantum circuit (VQC) models is the issue of barren plateaus leading to vanishing gradients. Classical Residual Networks (ResNet) have profoundly impacted deep learning by enabling the training of extremely deep neural networks through the introduction of skip connections that mitigate the vanishing gradient problem he2015deep ; marion2022scaling . This suggests that the implementation of Residual Networks within a quantum variational model could provide a significant advancement in quantum machine learning if they could similarly be utilised to avoid the problem of barren plateaus inherent in many VQC models. This line of reasoning was suggested in a recent review of barren plateaus larocca2024review with the caveat that the no-cloning theorem may make residual and skipped connections difficult to achieve in quantum circuits. In this work, we shall show a probabilistic implementation of a quantum native ResNet that does not require cloning states and could show promise against barren plateaus in VQC models.
@@
There has been active interest in building quantum ResNet type algorithms in the literature, with a primary focus being on quantum-classical hybrid models, that use the powerful modern architecture of classical ResNet models, while including quantum machine learning subroutines to produce a hybrid algorithm HASSAN2024105560 ; Sagingalieva_2023 ; zaman2024comparative . There have also been proposals to implement quantum residual connections by utilising several quantum neural network models in series with residual connections between them Kashif_2024 . A native VQC implementation of a quantum ResNet algorithm has previously been introduced in CrognalettiQResNet , in which skipped connections are possible in a VQC model if the implemented variational layers are restricted to be of the form of a unitary circuit followed by a product Pauli encoding, followed by the unitary circuit conjugate. This implements the quantum ResNet natively on a quantum device; however, as the authors note, this does not cover any general unitary variational layer $W_{l}(\mathbf{\theta}_{l})$, since implementing residual connections for a general $W_{l}(\mathbf{\theta}_{l})$ would not be a unitary process overall and was therefore not considered in this approach. A common theme explored in these previous works was utilising the effectiveness of classical ResNet models in tackling the vanishing gradients problem of deep neural networks and applying this reasoning in a quantum setting to the problem of barren plateaus.
@@
whereby a portion of the previous state $\lvert\psi_{l-1}\rangle$ is able to skip the variational operator $W_{l}(\mathbf{\theta}_{l})$ in that layer. This allows the overall model cost function to retain terms that have only passed through one variational layer, corresponding to very short circuits, while still providing terms that have been passed through all layers and hence potentially very deep circuits. The operator which would result in this would be of the form
@@
1.2 Average Pooling Layers
@@
Classical Convolution Neural Networks (CNN) are of significant importance to machine learning, mainly due to their structured manner of handling image and video data RawatCNN ; SHARMA2018377 ; SaffarCNN . Inspiration from these models has led to the development of quantum analogues. Quantum convolution neural networks have been previously proposed Cong_2019 ; umeano2023learn to classify quantum states with certain symmetry-protected topological phases and to classify image datasets Hur_2022 ; LiYao_2020 . In these models variational circuits are used during convolutional layers followed by further variational circuits and measurements in the pooling layer to perform dimensionality reduction, such that the operations are performed in a manner that respects certain symmetries of the data. Significant benefits of these models, such as avoiding barren plateaus Pesah_2021 , have been identified. Quantum-classical hybrid techniques have also been developed that utilise classical convolution neural network architectures alongside quantum models Liu_2021 .
@@
The implementation we present here differs in that we focus entirely on implementing a subroutine of classical CNN models, the average pooling layer, for amplitude encoded image data. We consider the subroutine where a pooling window of size $D\times D$ passes over the image and outputs the average of all pixels found within the pooling window. We show that this can be implemented natively on a quantum circuit by utilising the LCU method and further that this potentially leads to quadratic speedup in $N$, for images of dimension $N\times N$, over performing the subroutine’s classical implementation due to quantum parallelism. This demonstrates the possible utility of the LCU method, while providing a potentially advantageous subroutine for future quantum convolutional neural network models.
1.3 Irreducible Representation Subspace Symmetry Projections
@@
Figure 1: Visual representation of the irreducible subspace projection of a quantum state. The input state $\lvert\psi\rangle$ has components in each irreducible representation subspace denoted by $\lvert\psi_{r}\rangle$. a) The component of $\lvert\psi\rangle$ in the subspaces can be partially amplified to any desired ratio between the subspace components. Controlling the amplification of the permutation symmetric subspace $\lvert\psi_{r=1}\rangle$ for point cloud data is the focus of Section 4.4. The state can also be fully projected to $\lvert\psi_{r=1}\rangle$ which for the permutation group $S_{n}$ would correspond to a fully permutation invariant state. Permutation invariant encodings have been the subject of previous work heredge2023permutation and are reviewed within our new framework in Section 4.3.1. b) The input state $\lvert\psi\rangle$ is fully projected to $\lvert\psi_{r=3}\rangle$ within the $r=3$ irreducible subspace. In Section 4.3.2 we show that this corresponds to creating rotationally invariant encodings of point cloud data.
@@
A known issue in Quantum Machine Learning (QML) is that as the number of qubits increases there is a decrease in the generalisation performance of algorithms huang_power_2021 . A common empirical explanation of this is that the exponentially large Hilbert space leads to an overly expressive feature encoding where overtraining on the data becomes commonplace. Without an accompanied exponential increase in the training data, this leads to an overall reduction in the performance in the validation data set. A possible solution to the issue of overtraining is to reduce the expressibility of the encoding. Examples of techniques that have attempted this range from projecting kernels to a lower-dimensional space inductive_bias and approaches that are capable of encoding inductive biases directly into quantum states bowles2023contextuality . Furthermore, geometric QML techniques have studied methods for creating variational circuits that are equivariant with respect to data symmetries meyer22 ; Nguyen22 ; west2024provably ; Schatzki_2024 ; tüysüz2024symmetry in similar attempts to reduce the expressibility of QML models. We instead focus on implementing symmetries directly into the quantum encoded data, which is a stricter implementation of symmetry, meaning our procedure is agnostic to the trainable classification procedure.
@@
In previous work a quantum encoding was proposed using permutation symmetry, which led to a reduction in the dimensionality of the encoding and improved classification performance heredge2023permutation . This was discussed in the context of point cloud data (unordered collection of points in 3 dimensions that collectively represent an image), where each point cloud $X$ consists of $n$ points and each point $\textbf{p}_{i}$ is a 3-dimensional vector. As points do not have intrinsic ordering, the ordering of the points as they are input into a classification algorithm should ideally have no effect on the outcome. Therefore, the point cloud data naturally has point ordering permutation invariance. In general, a machine learning classifier $f$, could return a different result depending on the order of the points, that is, $f(\textbf{p}_{1},\textbf{p}_{2})\neq f(\textbf{p}_{2},\textbf{p}_{1})$, unless it has been specifically constructed to respect the permutation symmetry. The permutation invariant encoding heredge2023permutation acts by creating an equal quantum superposition of all permutations of the data. In the two-state case, given two points $\textbf{p}_{1},\textbf{p}_{2}$ encoded into the quantum state $\lvert\textbf{p}_{1}\rangle\otimes\lvert\textbf{p}_{2}\rangle$ this would correspond to preparing the permutation invariant state
@@
cannot be implemented via a unitary operation buzek_optimal_2000 . However, this process can be implemented in a probabilistic manner using ancilla qubits barenco1996 . This was shown to lead to improved classification performance for point cloud image classification heredge2023permutation .
@@
In this work, we demonstrate a generalisation of this technique that allows projections to any irreducible representation subspace of a finite group. Unlike previous works heredge2023permutation , this means that we are no longer restricted to the symmetric subspace or the permutation group $S_{n}$. Furthermore, our technique allows for linear combinations of projections to any irreducible representation subspaces.
@@
Utilising this framework, we demonstrate a rotationally invariant encoding for point cloud data as an example use case. This encoding produces the same quantum encoded state each time, even if the data input point cloud is rotated by any amount in 3-dimensions. We show that this is achieved both theoretically and numerically. This is a highly desirable property of the model as point cloud data naturally has rotational symmetry. Especially in applications such as computer vision for autonomous vehicles it is of upmost importance that subjects in the image, such as pedestrians, are correctly identified regardless of the angle from which they are being viewed. Previous work has focused on implementing rotational and permutation symmetry into models for point cloud data li2024enforcing , which did so by implementing an equivariant variational model. We highlight that our work does not use equivariant variational models, but projects quantum input states to a rotationally invariant subspace, hence strictly enforcing rotational invariance into any model for which this input encoded state is passed, as we effectively delete all information of the state which is not rotationally invariant.

### v2

@@ Abstract
We introduce several probabilistic quantum algorithms that overcome the normal unitary restrictions in quantum machine learning by leveraging the Linear Combination of Unitaries (LCU) method. Among our investigations are quantum native implementations of Residual Networks (ResNet), where we show that residual connections between layers of a variational ansatz can prevent barren plateaus in models which would otherwise contain them. Secondly, we implement a quantum analogue of average pooling layers from convolutional networks using single qubit controlled basic arithmetic operators and show that the LCU success probability remains stable for the MNIST database. This method can be further generalised to convolutional filters, while using exponentially fewer controlled unitaries than previous approaches. Finally, we propose a general framework for applying a linear combination of irreducible subspace projections on quantum encoded data. This enables a quantum state to remain within an exponentially large space, while selectively amplifying specific subspaces relative to others, alleviating simulability concerns that arise when fully projecting to a polynomially sized subspace. We demonstrate improved classification performance for partially amplified permutation invariant encoded point cloud data when compared to non-invariant or fully permutation invariant encodings. We also demonstrate a novel rotationally invariant encoding for point cloud data via Schur-Weyl duality. These quantum computing frameworks are all constructed using the LCU method, suggesting that further novel quantum machine learning algorithms could be created by utilising the LCU technique.
I Introduction
@@
Quantum computing is an emerging technology with the potential to solve certain problems far more efficiently than classical techniques Shor (1997); Cao et al. (2019); Herman et al. (2022). The search for useful applications of quantum computers within the domain of machine learning is of particular importance given the significant impact classical machine learning has had on many fields and industries in recent years Touvron et al. (2023); Chang et al. (2023); Frieder et al. (2023). Various quantum machine learning algorithms have been proposed, with some of the most common examples attempting to recreate a quantum analogue of the classical neural network Abbas et al. (2021); Beer et al. (2020). However, the inherently unitary operations of quantum algorithms impose significant limitations, as they restrict the implementation of non-unitary operations that are essential in many classical machine learning models. To address this issue, we explore the Linear Combination of Unitaries (LCU) method Andrew Childs (2012); Berry et al. (2015), a technique that employs ancilla qubits to probabilistically implement a linear combination of unitary operations, which therefore permits the implementation of non-unitary operations in quantum circuits. This unlocks the possibility of implementing non-unitary operations within quantum machine learning (QML) applications. We utilise this to present several novel applications of the LCU method with various advantages in the field of quantum machine learning.
@@
In Section II we adapt the classical residual learning framework to quantum variational circuits, facilitating a quantum native implementation of Residual Networks (ResNet) He et al. (2015). By partially skipping layers of the network, classical residual networks are able to avoid the vanishing gradient problem by allowing the gradients to flow through shallower sections of the network He et al. (2015); Marion et al. (2022). We show that a quantum ResNet may similarly provide a method of avoiding barren plateaus in Variational Quantum Circuit (VQC) models, by maintaining shallow depth contributions in the final loss function. Furthermore, we show that by including terms that parameterise the strength of the residual connections, it is possible to increase the lower bound of the probability of success of the LCU procedure helping to alleviate one of the LCU method’s principal issues.
@@
In Section III we implement a quantum analogue of average pooling operations from convolutional neural networks (CNN) Rawat and Wang (2017); Sharma et al. (2018); Al-Saffar et al. (2017) by using the LCU method. This provides an efficient implementation of average pooling for amplitude encoded image data which we demonstrate for any size pooling window. This may further be generalised to convolutional filters leading to an exponential improvement in the number of controlled unitaries required compared to previous techniques Wei et al. (2022); Chen et al. (2022). We demonstrate that the LCU probability of success will equal $1$ in the case where all pixels are of the same colour and will decrease when pixels in the local pooling window of dimension $D\times D$ become more diverse. For real-world images, we provide the intuition that the probability remains relatively stable since most pixels are similar locally, except at the edges of subjects in the image. This intuition is supported by empirical evidence on $N\times N$ pixel images in the MNIST Deng (2012) database, which shows that probability decreases but levels off to a finite value as $D$ increases, and shows no discernible trend when increasing image size $N$.
@@
In Section IV we present a method for projecting quantum encoded data to any combination of irreducible representation subspaces of a finite group, presenting a general framework for implementing full or partial symmetries in the encoding step of quantum machine learning models. These techniques are intended to reduce the effective dimension of quantum encoded data in an effort to improve generalisation performance, which has been reported to decline as the number of qubits and hence the dimension of the encoded quantum states increases Huang et al. (2021). We show that this technique can recreate previous work on permutation invariant encodings for point cloud data Heredge et al. (2024) as a special case. Furthermore, we implement a novel rotationally invariant encoding for point cloud data using the new technique by leveraging Schur-Weyl duality. This results in an encoded quantum state for point cloud data that is invariant if classical input point cloud data is rotated in 3-dimensional space, hence strongly enforcing rotation invariance on any model that subsequently uses this rotationally invariant input state. We further show that any combination of projections can be implemented at once, allowing certain symmetric subspaces of the data to be amplified or contracted to give increased flexibility over the amount of symmetry in the encoding. We demonstrate that intermediate levels of permutation symmetry for point cloud encoded data leads to an improved classification performance when compared to a non-symmetric or a fully permutation symmetric encoding.
@@
Table 1: Summary of the algorithms introduced in this work using the LCU framework. $G$ is a finite group. $\lvert\psi_{r}\rangle$ is the portion of a state $\lvert\psi\rangle$ that occupies the subspace of the irreducible representation $r$, with $\lvert\psi_{\text{rot}}\rangle$ meaning the portion of $\lvert\psi\rangle$ that is in the rotationally invariant subspace. Note that $\langle\psi_{r}\rvert\psi_{r}\rangle$ is not necessarily equal to 1, as the condition is $\langle\psi\rvert\psi\rangle=\sum_{r}\langle\psi_{r}\rvert\psi_{r}\rangle$ = 1. The parameters $a_{r}$ can be chosen by the user subject to normalisation constraints. In average pooling we use a pooling window of dimension $D\times D$ pixels. $L$ is the number of layers in the quantum ResNet. $\beta_{l}\in[0,1],\beta_{l}\in\mathbb{R}$ parametrise the strength of the skipped connections in the quantum ResNet. Throughout this work $\log\equiv\log_{2}$ is the logarithmic function for base 2.
Algorithm Result Ancilla Qubits Success Probability Section
@@
Quantum ResNet Residual Layers Residual connections can avoid barren plateaus $\mathcal{O}(L)$ $[\prod_{l}\big{(}1-4\beta_{l}(1-\beta_{l})\big{)},1]$ Sec II
Average Pooling Exponentially fewer controlled unitaries in averaging / convolutional step than previous techniques Wei et al. (2022); Chen et al. (2022) $\mathcal{O}(\log(D))$ Shown for MNIST in Fig 14 and Fig 15. Sec III
@@
Irreducible Subspace Projections Can freely enforce and parameterise symmetry of a quantum encoded state $\mathcal{O}(\log(\rvert G\rvert))$ $\sum_{r}\rvert a_{r}\rvert^{2}\langle\psi_{r}\rvert\psi_{r}\rangle$ Sec IV
@@
Rotationally Invariant Encoding Encoding respects rotational invariance of point cloud data $\mathcal{O}(\log(\rvert G\rvert))$ $\langle\psi_{\text{rot}}\rvert\psi_{\text{rot}}\rangle$ Sec IV.3.2
I.1 Quantum Native ResNet
@@
A common issue in the training of variational quantum circuit (VQC) models is the issue of barren plateaus leading to vanishing gradients. Classical Residual Networks (ResNet) have profoundly impacted deep learning by enabling the training of extremely deep neural networks through the introduction of skip connections that mitigate the vanishing gradient problem He et al. (2015); Marion et al. (2022). This suggests that the implementation of Residual Networks within a quantum variational model could provide a significant advancement in quantum machine learning if they could similarly be utilised to avoid the problem of barren plateaus inherent in many VQC models. This line of reasoning was suggested in a recent review of barren plateaus Larocca et al. (2024) with the caveat that the no-cloning theorem may make residual and skipped connections difficult to achieve in quantum circuits. In this work, we shall show a probabilistic implementation of a quantum native ResNet that does not require cloning states and could show promise against barren plateaus in VQC models.
@@
There has been active interest in building quantum ResNet inspired algorithms in the literature, with a primary focus being on quantum-classical hybrid models that use the powerful modern architecture of classical ResNet models, while including quantum machine learning subroutines to produce a hybrid algorithm Hassan et al. (2024); Sagingalieva et al. (2023); Zaman et al. (2024). There have also been proposals to implement quantum residual connections by utilising several quantum neural network models in series with residual connections between them Kashif and Al-Kuwari (2024). A native VQC implementation of a quantum ResNet algorithm has previously been introduced in Crognaletti et al. (2024, ), in which skipped connections are possible in a VQC model if the implemented variational layers are restricted to be of the form of a unitary circuit followed by a product Pauli encoding, followed by the unitary circuit conjugate. This implements the quantum ResNet natively on a quantum device; however, as the authors note, this does not cover any general unitary variational layer $W_{l}(\mathbf{\theta}_{l})$, since implementing residual connections for a general $W_{l}(\mathbf{\theta}_{l})$ would not be a unitary process overall and was therefore not considered in this approach. A common theme explored in these previous works was utilising the effectiveness of classical ResNet models in tackling the vanishing gradients problem of deep neural networks and applying this reasoning in a quantum setting to the problem of barren plateaus.
@@
whereby a portion of the previous state $\lvert\psi_{l-1}\rangle$ is able to skip the variational operator $W_{l}(\mathbf{\theta}_{l})$ in that layer. This allows the overall model cost function to retain terms that have only passed through one variational layer, corresponding to very shallow circuits, while still providing terms that have been passed through all layers and hence potentially very deep circuits. The operator which would result in this would be of the form
@@
The use of ancilla qubits to implement ResNet-like architectures in QML models has previously been explored Wen et al. (2024), revealing that residual connections within the data encoding segment of a circuit can expand the frequency spectrum of the resulting model, leading to more expressive encodings. In contrast, our work introduces a framework for quantum ResNets exclusively within the variational portion of a VQC model. We provide a detailed proof of the probability of success of the LCU procedure, demonstrating that the lower bound of this probability can be adjusted by varying the residual connection strength. Additionally, we illustrate that applying quantum residual layers to a model can mitigate the occurrence of barren plateaus in circuits which would otherwise contain them. We also demonstrate that quantum ResNets can be viewed as equivalent to ensembles of unitary VQC models with additional non-unitary terms. While we show that quantum ResNets can avoid barren plateaus, we also discuss how they may likely be classically simulatable in many cases, at least in the case that the connection between absence of barren plateaus and classical simulatability Cerezo et al. (2023) is valid for the constituent components. We propose the solution to this quantum ResNet simulatability issue may lie in the non-unitary terms and suggest a characterisation of these terms as a topic for further research.
I.2 Average Pooling Layers
@@
Classical Convolution Neural Networks (CNN) are of significant importance to machine learning, mainly due to their structured manner of handling image and video data Rawat and Wang (2017); Sharma et al. (2018); Al-Saffar et al. (2017); Vu et al. (2024). Inspiration from these models has led to the development of quantum analogues. Quantum convolution neural networks have been previously proposed Cong et al. (2019); Umeano et al. (2023) to classify quantum states with certain symmetry-protected topological phases and to classify image datasets Hur et al. (2022); Li et al. (2020). In these models variational circuits are used during convolutional layers followed by further variational circuits and measurements in the pooling layer to perform dimensionality reduction such that the operations are performed in a manner that respects certain symmetries of the data. Significant benefits of these models, such as avoiding barren plateaus Pesah et al. (2021), have been identified. Quantum-classical hybrid techniques have also been developed that utilise classical convolution neural network architectures alongside quantum models Liu et al. (2021).
@@
The implementation we present here differs in that we focus entirely on implementing a subroutine of classical CNN models, the average pooling layer, for amplitude encoded image data. We consider the subroutine where a pooling window of size $D\times D$ passes over the image and outputs the average of all pixels found within the pooling window. We show that this can be implemented natively on a quantum circuit by utilising the LCU method which could lead to an improvement over performing the subroutine classically, as quantum parallelism allows the averaging operation to apply to all pixel simultaneously. This demonstrates the possible utility of the LCU method, while providing a potentially advantageous subroutine for future quantum convolutional neural network models.
Previous work has investigated the LCU technique for creating convolutional layer filters Wei et al. (2022) based on spatial filtering Yao et al. (2017), which can recover average pooling as a special case. This work did not generalise to any size $D$, restricting instead to $D=3$, but stated the method would be require $D^{2}$ multi-controlled operators in general. In contrast, we show a valid proof for any $D$ and an efficient circuit implementation which requires only $\mathcal{O}(\log(D))$ single-qubit controlled unitaries, leading to an exponential improvement in $D$ over previous work Wei et al. (2022). We also show that our construction can be generalised by adjusting the ancilla qubit state initialisation in order to implement a general convolutional layer filter, recovering the main result of Wei et al. (2022) while maintaining an exponential improvement in $D$.
I.3 Irreducible Representation Subspace Symmetry Projections
@@
Figure 1: Visual representation of the irreducible subspace projection of a quantum state. The input state $\lvert\psi\rangle$ has components in each irreducible representation subspace denoted by $\lvert\psi_{r}\rangle$. a) The component of $\lvert\psi\rangle$ in the subspaces can be partially amplified to any desired ratio between the subspace components. Controlling the amplification of the permutation symmetric subspace $\lvert\psi_{r=1}\rangle$ for point cloud data is the focus of Section IV.4. The state can also be fully projected to $\lvert\psi_{r=1}\rangle$ which for the permutation group $S_{n}$ would correspond to a fully permutation invariant state. Permutation invariant encodings have been the subject of previous work Heredge et al. (2024) and are reviewed within our new framework in Section IV.3.1. b) The input state $\lvert\psi\rangle$ is fully projected to $\lvert\psi_{r=3}\rangle$ within the $r=3$ irreducible subspace. In Section IV.3.2 we show that this corresponds to creating rotationally invariant encodings of point cloud data.
@@
A known issue in Quantum Machine Learning (QML) is that as the number of qubits increases there is a decrease in the generalisation performance of algorithms Huang et al. (2021). A common empirical explanation of this is that the exponentially large Hilbert space leads to an overly expressive feature encoding where overtraining on the data becomes commonplace. Without an accompanied exponential increase in the training data, this leads to an overall reduction in the performance in the validation data set. A possible solution to the issue of overtraining is to reduce the expressibility of the encoding. Examples of techniques that have attempted this range from projecting kernels to a lower-dimensional space Kübler et al. (2021) and approaches that are capable of encoding inductive biases directly into quantum states Bowles et al. (2023). Furthermore, geometric QML techniques have studied methods for creating variational circuits that are equivariant with respect to data symmetries Meyer et al. (2022); Nguyen et al. (2022); West et al. (2024); Schatzki et al. (2024); Tüysüz et al. (2024) in similar attempts to reduce the expressibility of QML models. We instead focus on implementing symmetries directly into the quantum encoded data, which is a stricter implementation of symmetry, meaning our procedure is agnostic to the trainable classification procedure.
@@
In previous work a quantum encoding was proposed using permutation symmetry, which led to a reduction in the dimensionality of the encoding and improved classification performance Heredge et al. (2024). This was discussed in the context of point cloud data (unordered collection of points in 3 dimensions that collectively represent an image), where each point cloud $X$ consists of $n$ points and each point $\textbf{p}_{i}$ is a 3-dimensional vector. As points do not have intrinsic ordering, the ordering of the points as they are input into a classification algorithm should ideally have no effect on the outcome. Therefore, the point cloud data naturally has point ordering permutation invariance. In general, a machine learning classifier $f$, could return a different result depending on the order of the points, that is, $f(\textbf{p}_{1},\textbf{p}_{2})\neq f(\textbf{p}_{2},\textbf{p}_{1})$, unless it has been specifically constructed to respect the permutation symmetry. The permutation invariant encoding Heredge et al. (2024) acts by creating an equal quantum superposition of all permutations of the data. In the two-state case, given two points $\textbf{p}_{1},\textbf{p}_{2}$ encoded into the quantum state $\lvert\textbf{p}_{1}\rangle\otimes\lvert\textbf{p}_{2}\rangle$ this would correspond to preparing the permutation invariant state
@@
cannot be implemented via a unitary operation Buzek and Hillery (2000). However, this process can be implemented in a probabilistic manner using ancilla qubits Barenco et al. (1996). This was shown to lead to improved classification performance for point cloud image classification Heredge et al. (2024).
@@
In this work, we demonstrate a generalisation of this technique that allows projections to any irreducible representation subspace of a finite group. Unlike previous works Heredge et al. (2024), this means that we are no longer restricted to the symmetric subspace or the permutation group $S_{n}$. Furthermore, our technique allows for linear combinations of projections to any irreducible representation subspaces.
@@
Utilising this framework, we demonstrate a rotationally invariant encoding for point cloud data as an example use case. This encoding produces the same quantum encoded state each time, even if the data input point cloud is rotated by any amount in 3-dimensions. We show that this is achieved both theoretically and numerically. This is a highly desirable property of the model as point cloud data naturally has rotational symmetry. Especially in applications such as computer vision for autonomous vehicles it is of upmost importance that subjects in the image, such as pedestrians, are correctly identified regardless of the angle from which they are being viewed. Previous work has focused on implementing rotational and permutation symmetry in models for point cloud data Li et al. (2024), which did so by implementing an equivariant variational model. We highlight that our work does not use equivariant variational models but projects quantum input states to a rotationally invariant subspace, hence strictly enforcing rotational invariance into any model for which this input encoded state is passed, as we effectively delete all information of the state which is not rotationally invariant.